what voulme of 0.550 M Na2S are needed to react with 15.00mL of 0.217 M Agno3


1
Expert's answer
2020-08-17T13:34:21-0400

LET VOLUME IS V ml .


Moles of Na2S = V * 0.550 / 1000 .


Moles Of AgNO3 = 15 * 0.217 / 1000.


Moles Of AgNO3 = 2 * V * 0.550 / 1000 = 15 * 0.217 / 1000 .


So , V = 15 * 0.217 / 2 * 0.550 = 2.9590 ml ...........